OpenStudy (anonymous):

Solution set for (x^2-6x)^2-35(x^2-6x)-200=0

myininaya (myininaya):

let u=x^2-6x then u^2=(x^2-6x)^2 so we have u^2-35u-200=0 \[u=\frac{35 \pm \sqrt{(-35)^2-4(1)(-200)}}{2(1)}=\frac{35 \pm \sqrt{1225+800}}{2}=\frac{35 \pm \sqrt{2025}}{2}\] \[=\frac{35 \pm 45}{2}=\frac{35+45}{2},\frac{35-45}{2}=\frac{80}{2}.\frac{-10}{2}=40,-5\] but remember u=x^2-6x so x^2-6x=40 or x^2-6x=-5 solve both of these equations for x